# Scaling of Disorder Operator and Entanglement Entropy at Easy-Plane Deconfined Quantum Criticalities

### v1

@@ Abstract
We systematically investigate the scaling behavior of the disorder operator and the entanglement entropy (EE) of the easy-plane JQ (EPJQ) model at its transition between the antiferromagnetic XY ordered phase (AFXY) and the valence bond solid (VBS) phase. We find (1) there exists a finite value of the order parameters at the AFXY-VBS phase transition points of the EPJQ model, and the finite order parameter is strengthened as anisotropy $\Delta$ varies from the Heisenberg limit ($\Delta=1$) to the easy-plane limit ($\Delta=0$); (2) Both EE and disorder operator with smooth boundary cut exhibit anomalous scaling behavior at the transition points, resembling the scaling inside the Goldstone model phase, and the anomalous scaling becomes strengthened as the transition becomes more first order; (3) First put forward in Ref. [arXiv:2401.12838], with the finite-size corrections in EE for Goldstone phase is properly considered in the fitting form, the anomalous scaling behavior of EE can be adapted with emergent SO(5) symmetry breaking at the Heisenberg limit ($\Delta=1$). We extend this method in the EPJQ model and observe similar results, which may indicate emergent SO(4) symmetry breaking in the easy-plane regime ($\Delta<1$) or emergent SO(5) symmetry breaking in the Heisenberg limit ($\Delta=1$). These observations provide evidence that the Néel-VBS transition in the JQ model setting evolves from weak to prominent first-order transition as the system becomes anisotropic, and the non-local probes such as EE and disorder operator, serve as the sensitive tool to detect such salient yet fundamental features.
@@ I Introduction
The conventional Landau-Ginzburg-Wilson (LGW) paradigm for characterizing phases and their transitions has been confronted with the notion of deconfined quantum critical points (DQCPs), which describes a direct continuous transition between Néel ordered phases and valence-bond-solid (VBS) phases Senthil (2004); Senthil et al. (2004); Levin and Senthil (2004); Senthil et al. (2005). However, the feasibility of realizing such transitions in lattice models remains a contentious issue because of the observation of anomalous behavior against conformal field theories (CFT) Harada et al. (2013); Chen et al. (2013); Nahum et al. (2015); Nakayama and Ohtsuki (2016); Li (2022); Poland et al. (2019); Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024), such as drifting of critical exponents Harada et al. (2013); Chen et al. (2013); Nahum et al. (2015), possible violation of CFT bounds Nakayama and Ohtsuki (2016); Li (2022); Poland et al. (2019), and the anomalous scaling behavior of entanglement entropies and disorder operators Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024). A deeper exploration of the origins of these anomalous behaviors will benefit not only the better understanding of nature of DQCPs but also its lattice model and even material realizations Zayed et al. (2017); Guo et al. (2020); Jiménez et al. (2021); Sun et al. (2021); Cui et al. (2023); Guo et al. (2023).
@@
Among all the anomalous behavior against CFTs, we are in particular interested in the anomalous finite-size scaling behavior of Rényi EE and disorder operator, as it happens even at relatively small system sizes Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024) where other physical quantities such as order parameters and correlation functions still exhibit good agreement with a continuous phase transition Sandvik (2007); Lou et al. (2009); Shao et al. (2016); Ma et al. (2018a); Qin et al. (2017). In fact, at the DQCPs of SU(2) JQ2 and JQ3 models, our previous works Zhao et al. (2022a); Song et al. (2023a); Wang et al. (2022); Song et al. (2023b) show that disorder operator and EE all exhibit anomalous scaling against CFTs. To be specific, for the corner cut case where the boundary of subregion has four $\frac{\pi}{2}$ corners, both disorder operator and EE have a positive logarithmic correction to the leading order area law term which violates the unitary CFT prediction that the correction must be negative  Casini and Huerta (2012); Fradkin and Moore (2006); Casini and Huerta (2007). Later, further studies clarified that the violation of EE actually arises from the smooth part of the subregion. In fact, for smooth cut case where the boundary has no sharp corners, the anomalous logarithmic subleading correction also exists in EE, which is incompatible with CFTs Song et al. (2023b, a), but seems to be more compatible with the existence of Goldstone modes  Metlitski and Grover (2015); Deng et al. (2024) and can also possibly be adapted with the ‘walking" pseudo-criticality behavior at the transition Wang et al. (2017); Nahum (2020); Ma and Wang (2020); Zhou et al. (2023). In addition, recently another work D’Emidio and Sandvik (2024) shows that scaling of EE at the DQCP is cut-dependent, and the anomalous log correction for smooth cut can be suppressed by considering a tilted square lattice by $\frac{\pi}{4}$ degrees, and the essential information of the CFTs might be captured by this cut at small system sizes. In this context, the scalings of these nonlocal physical observables prove to be powerful and sensitive tools for diagnosing the various possible scenarios of the DQCPs Zhao et al. (2022a); Song et al. (2023a); Wang et al. (2022); Song et al. (2023b); Deng et al. (2024) and revealing the universal information of the possible CFT nearby the transitions D’Emidio and Sandvik (2024).
@@
In this paper, we use the stochastic series expansion (SSE) quantum Monte Carlo (QMC) methods Sandvik (1999); Syljuåsen and Sandvik (2002) to systematically study the order parameter, entanglement entropy, and disorder operator at the transitions of the EPJQ model. We carefully analyze the remaining ordered moments at the transitions and systematically study the evolution of the scaling of entanglement entropy and disorder operator along the transition line, as a function of $\Delta$ in Eq. (1). Our major findings are
@@
    For $\Delta\in[0,1)$ there always exist finite value of the order parameters at the Néel-VBS phase transitions of the model (including the finite order parameter at the $\Delta=1$ limit Takahashi et al. (2024)), indicating the first-order nature of the transition, and the moment increases as $\Delta$ is tuned from 1 to 0;
@@
    Both EE and disorder operator with smooth boundary cut exhibit anomalous scaling behavior at the transition points, resembling the scaling inside the Goldstone model phase, and the anomalous scaling becomes strengthened as the transition becomes more first order;
@@
    Pioneered in Ref. Deng et al. (2023a, 2024), with the finite-size corrections in EE for Goldstone phase is properly dealt with, the anomalous scaling behavior of EE can be adapted with emergent SO(5) symmetry breaking at the Heisenberg limit ($\Delta=1$). We extend this method in EPJQ model and observe similar but weaker results which may indicate emergent SO(4) symmetry breaking in the easy-plane regime ($\Delta<1$) or emergent SO(5) symmetry breaking in the Heisenberg limit ($\Delta=1$).
Our work thus provides a comprehensive study of the scaling of EE and disorder operator along varying $\Delta$ from 1 to 0 where the weakly first-order transition is gradually tuned to a prominent first-order one. We show that the anomalous scaling of EE and disorder operator is related with the first order behavior, proving a strong evidence that the DQCP of JQ model is indeed a weakly-first-order transition and the anomalous log-corrections are from Goldstone modes. Apart from this, our work supports the idea in Ref. Deng et al. (2024) that the origin of anomalous scaling of EE at the transition of $\Delta=1$ is from the a continuous SO(5) symmetry breaking and the Goldstone modes associated with it. However, we do not observe the same good convergence behavior to exact number of Goldstone modes required by the SO(5) symmetry breaking at $\Delta=1$ or SO(4) symmetry breaking at $\Delta<1$, which might originate from insufficient data quality and the less pronounced emergent SO(4) symmetry in the easy-plane case. Despite that, our results strongly suggest that the DQCPs of EPJQ model is first-order and scaling of EE and disorder operator serve as sensitive tool for detecting such transitions.
The rest of the paper is organized as follows. In Sec. II, we first introduce the EPJQ model and discuss how we extract the finite order parameters at its transition points, then we move on to the discussion of disorder operator in Sec. III and reveal its anomalous scaling in comparison with that of the conventional (2+1)D O(3) QCP and Néel phase, next we discuss the results of EE at the EPJQ transitions in Sec. IV and point out its anomalous scaling which might stem from the Goldstone mode at the first-order transition points, finally we give a comprehensive summary on our results, discuss the relation of our work with other recent works Deng et al. (2024); D’Emidio and Sandvik (2024) and point out future directions in Sec. V. Detailed analysis of the determination of the transition points and extrapolations of the order parameter, and the analysis of the quality of the fitting in EE data are presented in the Supplemental Material (SM) sup .
@@ II Model and Phase Diagram
We study the easy-plane JQ3 (EPJQ) model Qin et al. (2017); Ma et al. (2019); Zhao et al. (2019); Desai and Kaul (2020); Sun et al. (2021) as illustrated in the left inset of Fig. 1 (a) with the following Hamiltonian
@@
where $P_{ij}=\frac{1}{4}-\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ is the two-spin singlet projector. For $\Delta=1$, the model goes back to JQ3 with isotropic Heisenberg interactions. It was found the quantum phase transition in the JQ3 model between the Néel and VBS phases at $q_{c}=0.59864(5)$ (with $q=\frac{Q}{J+Q}$ and $J+Q=1$ as the energy unit) Lou et al. (2009); Wang et al. (2022), but more numerical evidence, especially from the entanglement witnesses such as EE Zhao et al. (2022a); Song et al. (2023a, b) and disorder operator Wang et al. (2022), reveal an anomalous subleading log-correction even if the partitioning of the lattice is smooth, i.e. no sharp corners on the boundary. If the transition is indeed continuous and acquires a unitary conformal field theory (CFT) description, one shall not see the log-correction with smooth boundary Song et al. (2023a). The fact that the sign of the observed log-coefficient is consistent with that of the presence of the Goldstone mode, further suggests there exist finite antiferromagnetic moment, i.e. remaining of the Néel order, at the transition point Metlitski and Grover (2015). Such evidence promotes the understanding that the Néel-VBS transition is first order, despite being very weak at the SU(2) limit Takahashi et al. (2024). In this work, we will monitor the behavior of the transition as the anisotropy in Eq. (1) is tuned away from the Heisenberg limit.
@@
At each $\Delta\in[0,1)$, the phase transition at zero temperature happens between the antiferromagnetic XY (AFXY) and VBS phase when tunning $q=Q/(J+Q)$ from zero to $q_{c}$. We focus on the criticality and perform SSE-QMC simulations on the EPJQ model at $\beta=2L$ for different system sizes $L=16,24,32,48,64,96,128$ with $\Delta=0$, $0.1$, $0.3$, $0.5$ and $0.6$. In the AFXY phase, the order parameter should be the sublattice magnetization which breaks the U(1) symmetry, whose $x$ component is defined as
$\begin{split}M_{x}=\frac{1}{N}\sum_{i=1}^{N}(-1)^{x_{i}+y_{i}}S_{i}^{x}\end{split}$ (2)
with $\{x_{i}$,$y_{i}\}$ the coordinates of site $i$ in the lattice. In the simulation with finite system sizes the expectation value of $\langle M_{x}^{2}\rangle$ can be viewed as the square of the AFXY order parameter, whereas $\langle M_{x}\rangle=0$ for finite size.
For the VBS phase, the valence bond can form in horizontal and vertical directions as exemplified of the former in the right inset of Fig. 1 (a), which can be described through observable
@@
We thus perform the finite-size scaling (FSS) analysis of $\langle M^{2}_{x}\rangle$ and $\langle D^{2}\rangle$ to determine the transition points at the thermodynamic limit (TDL) Ma et al. (2018b). The dimensionless quantity Binder cumulants are widely used in the FSS. For those two ground states in our study, the Binder cumulants are defined as
@@
We have performed the extrapolation of the ordered moments as in Fig. 2 (d) for all the $\Delta$ values, and the obtained results are summarized in Fig. 1 (b). From which, one clearly sees that $\langle M^{2}_{x}\rangle$ and $\langle D^{2}\rangle$ are small at $\Delta=0.6$ (they are even smaller at the isotropic limit), but they gradually increases as $\Delta\to 0$. We show all these extrapolations in the Sec. I of SM sup . The enhancement of the ordered moments as $\Delta\to 0$ is the reason we believe along the phase AFXY-VBS phase boundary in Fig. 1 (a), the transition evolves from weakly first-order to stronger ones. As will be shown below, our non-local EE and disorder operator measurements (with smaller systems sizes than those of order parameters) successfully capture such understanding from the more traditional local measurements.
For comparison, we have also computed the disorder operator for the square lattice $J_{1}$-$J_{2}$ Heisenberg model with the Hamiltonian
@@
where $\langle ij\rangle$ denotes the $J_{1}$ bond and $\langle ij\rangle^{{}^{\prime}}$ denotes the $J_{2}$ bond. This model has a well-established (2+1)d O(3) QCP at $(J_{2}/J_{1})_{c}=1.90951(1)$ Ma et al. (2018b) separating the Néel phase and a symmetric singlet product phase. As will also shown below, the entanglement measurements (both disorder operator and EE) of this O(3) QCP do not have anomalous log-corrections.

### v2

@@ Abstract
We systematically investigate the scaling behavior of the disorder operator and the entanglement entropy (EE) of the easy-plane JQ (EPJQ) model at its transitions between the antiferromagnetic XY ordered phase (AFXY) and the valence bond solid (VBS) phase. We find (1) there exists a tiny yet finite value of the order parameters at the AFXY-VBS phase transition points of the EPJQ model, and the finite order parameter is strengthened as anisotropy $\Delta$ varies from the Heisenberg limit ($\Delta=1$) to the easy-plane limit ($\Delta=0$); (2) Both EE and disorder operator with smooth boundary cut exhibit anomalous scaling behavior at the transition points, resembling the scaling inside the Goldstone model phase, and the anomalous scaling becomes strengthened as the transition becomes more first order; (3) First put forward in Ref. [arXiv:2401.12838], with the finite-size corrections in EE for Goldstone phase is properly considered in the fitting form, the anomalous scaling behavior of EE can be adapted with emergent SO(5) symmetry breaking at the Heisenberg limit ($\Delta=1$). We extend this method in the EPJQ model and observe similar yet weaker results, which may indicate emergent SO(4) symmetry breaking in the easy-plane regime ($\Delta<1$) or emergent SO(5) symmetry breaking in the Heisenberg limit ($\Delta=1$). These observations provide evidence that the Néel-VBS transition in the JQ model setting evolves from weak to prominent first-order transition as the system becomes anisotropic, and the non-local probes such as EE and disorder operator, serve as the sensitive tool to detect such salient yet fundamental features.
@@ I Introduction
The conventional Landau-Ginzburg-Wilson (LGW) paradigm for characterizing phases and their transitions has been confronted with the notion of deconfined quantum critical points (DQCPs), which describes a direct continuous transition between Néel ordered phases and valence-bond-solid (VBS) phases Senthil (2004); Senthil et al. (2004); Levin and Senthil (2004); Senthil et al. (2005). However, the feasibility of realizing such transitions in lattice models remains a contentious issue because of the observations of various anomalous behavior against conformal field theories (CFT) Harada et al. (2013); Chen et al. (2013); Nahum et al. (2015); Nakayama and Ohtsuki (2016); Li (2022); Poland et al. (2019); Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024), such as drifting of critical exponents Harada et al. (2013); Chen et al. (2013); Nahum et al. (2015), violation of CFT bounds Nakayama and Ohtsuki (2016); Li (2022); Poland et al. (2019), and the anomalous scaling behavior of entanglement entropies and disorder operators Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024). A deeper exploration of the origins of these anomalous behaviors will benefit not only the better understanding of nature of DQCPs but also its lattice model and even material realizations Zayed et al. (2017); Guo et al. (2020); Jiménez et al. (2021); Sun et al. (2021); Cui et al. (2023); Guo et al. (2023).
@@
Among all the anomalous behavior against CFTs, we are in particular interested in the anomalous finite-size scaling behavior of Rényi entanglement entropy (EE) and disorder operator, as it happens even at relatively small system sizes Wang et al. (2022); Zhao et al. (2022a); Song et al. (2023a, b); Liu et al. (2023); Liao et al. (2023); Liu et al. (2024) where other physical quantities such as order parameters and correlation functions still exhibit good agreement with a continuous phase transition Sandvik (2007); Lou et al. (2009); Shao et al. (2016); Ma et al. (2018a); Qin et al. (2017). In fact, at the DQCPs of SU(2) JQ2 and JQ3 models, our previous works Zhao et al. (2022a); Song et al. (2023a); Wang et al. (2022); Song et al. (2023b) show that disorder operator and EE all exhibit anomalous scaling against CFTs. To be specific, for the corner cut case where the boundary of subregion has four $\frac{\pi}{2}$ corners, both disorder operator and EE have a positive logarithmic correction to the leading order area law term which violates the unitary CFT prediction that the correction must be negative  Casini and Huerta (2012); Fradkin and Moore (2006); Casini and Huerta (2007). Later, further studies clarified that the violation of EE actually arises from the smooth part of the subregion. In fact, for smooth cut case where the boundary has no sharp corners, the anomalous logarithmic subleading correction also exists in EE, which is incompatible with CFTs Song et al. (2023b, a), but seems to be more compatible with the existence of Goldstone modes  Metlitski and Grover (2015); Deng et al. (2024) and can also possibly be adapted with the ‘walking" pseudo-criticality behavior at the transition Wang et al. (2017); Nahum (2020); Ma and Wang (2020); Zhou et al. (2023). In addition, recently another work D’Emidio and Sandvik (2024) shows that scaling of EE at the DQCP is cut-dependent, and the anomalous log correction for smooth cut can be suppressed by considering a tilted square lattice by $\frac{\pi}{4}$ degrees, and the essential information of the CFTs might be captured by this cut at small system sizes. In this context, the scalings of these nonlocal physical observables prove to be powerful and sensitive tools for diagnosing the various possible scenarios of the DQCPs Zhao et al. (2022a); Song et al. (2023a); Wang et al. (2022); Song et al. (2023b); Deng et al. (2024) and revealing the universal information of the possible CFT nearby the transitions D’Emidio and Sandvik (2024).
@@
In this paper, we use the stochastic series expansion (SSE) quantum Monte Carlo (QMC) methods Sandvik (1999); Syljuåsen and Sandvik (2002) to systematically study the order parameter, entanglement entropy, and disorder operator at the transitions of the EPJQ model. We carefully analyze the remaining ordered moments at the transitions and systematically study the evolution of the scaling of entanglement entropy and disorder operator along the critical line as a function of $\Delta$ in Eq. (1). Our major findings are
@@
    For $\Delta\in[0,1)$ there always exist tiny yet finite order parameters at the AFMXY/Néel-VBS phase transitions of the model (including the finite order parameter at the $\Delta=1$ limit Takahashi et al. (2024)), indicating the first-order nature of the transition, and the moment increases as $\Delta$ is tuned from 1 to 0;
@@
    Both EE and disorder operator with standard smooth boundary cut (without tilting) exhibit anomalous scaling behavior against CFTs at the transition points, and resemble more to the scalings inside the Goldstone model phase, and the anomalous scalings also become strengthened as the transition becomes more first order;
@@
    Pioneered in Ref. Deng et al. (2023a, 2024), with the finite-size corrections in EE for Goldstone phase is properly dealt with, the anomalous scaling behavior of EE can be adapted with emergent SO(5) symmetry breaking at the Heisenberg limit ($\Delta=1$). We extend this method in EPJQ model and observe similar yet weaker results which may indicate emergent SO(4) symmetry breaking in the easy-plane regime ($\Delta<1$) or emergent SO(5) symmetry breaking in the Heisenberg limit ($\Delta=1$).
Our work thus provides a comprehensive study of the scaling of EE and disorder operator along the critical line where the weakly first-order transition is gradually tuned to a prominent first-order one by tunning $\Delta=1$ to $\Delta=0$. We show that the anomalous scalings of EE and disorder operator are positively related with the first order behavior, proving a strong evidence that the DQCP of JQ model is indeed a very weakly-first-order transition and the anomalous log-corrections come from Goldstone modes. Apart from this, our work supports the findings in Ref. Deng et al. (2024) that the origin of anomalous scaling of EE at the transition of $\Delta=1$ is the emergent continuous SO(5) symmetry breaking and the Goldstone modes associated with it. However, our results are weaker as we do not observe the same good convergence behavior to exact number of Goldstone modes required by the SO(5) symmetry breaking at $\Delta=1$ or SO(4) symmetry breaking at $\Delta<1$, which might originate from insufficient data quality or the less pronounced emergent SO(4) symmetry behavior in the easy-plane case. Despite that, our results strongly suggest that the DQCPs of EPJQ model are first-ordered for all $\Delta$ and scalings of EE and disorder operator serve as sensitive tools for detecting such transitions.
The rest of the paper is organized as follows. In Sec. II, we first introduce the EPJQ model and discuss how we extract the finite order parameters at its transition points, then we move on to the discussion of disorder operator in Sec. III and reveal its anomalous scaling with comparisons to that of the conventional (2+1)D O(3) QCP and Néel phase, next we discuss the results of EE at the EPJQ transitions in Sec. IV and point out its anomalous scaling very likely stems from the residual Goldstone mode at the first-order transition points, finally we give a comprehensive summary on our results, discuss the relation of our work with other recent works Deng et al. (2024); D’Emidio and Sandvik (2024) and point out future directions in Sec. V. Detailed analysis of the determination of the transition points and extrapolations of the order parameter, and the analysis of the quality of the fitting in EE data are presented in the Supplemental Material (SM) sup .
@@ II Model and Phase Diagram
We study the easy-plane JQ3 (EPJQ) model Qin et al. (2017); Ma et al. (2019); Zhao et al. (2019); Desai and Kaul (2020); Sun et al. (2021) as illustrated in the lower left inset of Fig. 1 (a) with the following Hamiltonian
@@
where $P_{ij}=\frac{1}{4}-\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ is the two-spin singlet projector. For $\Delta=1$, the model reduces to JQ3 model with isotropic Heisenberg interactions. It has been found a direct quantum phase transition in the JQ3 model between the Néel and VBS phases happens at $q_{c}=0.59864(5)$ (with $q=\frac{Q}{J+Q}$ and $J+Q=1$ as the energy unit) Lou et al. (2009); Wang et al. (2022). However, more numerical evidence, especially from the scaling of nonlocal observables such as EE Zhao et al. (2022a); Song et al. (2023a, b) and disorder operator Wang et al. (2022), reveal an anomalous behavior against CFTs even if the partitioning of the lattice is smooth, i.e. no sharp corners on the boundary. The fact that the sign of the observed log-coefficient is consistent with that of the presence of the Goldstone mode, further suggests there exist finite antiferromagnetic moment, i.e. remaining of the Néel order, at the transition point Metlitski and Grover (2015). Such evidence promotes the understanding that the Néel-VBS transition is indeed first order, despite being very weak at the SU(2) limit Takahashi et al. (2024). In this work, we monitor the behavior of both conventional order parameters and nonlocal observables at such transitions, as the anisotropy $\Delta$ in Eq. (1) is tuned from the Heisenberg limit($\Delta=1$) to easy-plane limit ($\Delta=0$).
@@
At each $\Delta\in[0,1)$, the phase transition between the antiferromagnetic XY (AFXY) and VBS phase at zero temperature happens when tunning $q=Q/(J+Q)$ from zero to $q_{c}$. To determine the critical points, we perform SSE-QMC simulations Sandvik (1999); Syljuåsen and Sandvik (2002) on the EPJQ model at $\beta=2L$ for different system sizes $L=16,24,32,48,64,96,128$ with $\Delta=0$, $0.1$, $0.3$, $0.5$ and $0.6$. In the AFXY phase, the order parameter is defined as the sublattice magnetization which breaks the U(1) symmetry, with the $x$ component being written as
$\begin{split}M_{x}=\frac{1}{N}\sum_{i=1}^{N}(-1)^{x_{i}+y_{i}}S_{i}^{x},\end{split}$ (2)
where $\{x_{i}$,$y_{i}\}$ are coordinates of site $i$ in the lattice. In the simulation with finite system sizes, the expectation value of $\langle M_{x}^{2}\rangle$ can be viewed as the square of the AFXY order parameter, whereas $\langle M_{x}\rangle=0$ at finite sizes.
For the VBS phase, the valence bonds can form in horizontal or vertical directions as exemplified of the former in the right inset of Fig. 1 (a). This order can be quantified through observables
@@
We thus perform the finite-size scaling (FSS) analysis of $\langle M^{2}_{x}\rangle$ and $\langle D^{2}\rangle$ to determine the transition points at the thermodynamic limit (TDL) Ma et al. (2018b). The dimensionless quantity Binder cumulants are widely used in the FSS to extract the critical points. For those two ground states in our study, the Binder cumulants are defined as
@@
We have performed the extrapolation of the ordered moments as in Fig. 2 (d) for all the $\Delta$ values, and the obtained results are summarized in Fig. 1 (b). It is clear that $\langle M^{2}_{x}\rangle$ and $\langle D^{2}\rangle$ are small at $\Delta=0.6$ (they are even smaller at the isotropic limit), but they gradually increases as $\Delta\to 0$. We show all these extrapolations in the Sec. I of SM sup . The enhancement of the ordered moments as $\Delta\to 0$ along the phase AFXY-VBS phase boundary in Fig. 1 (a) suggests evolution from very weakly first-order to stronger ones. As will be shown below, our non-local EE and disorder operator measurements (with smaller systems sizes compared with order parameters) can also capture such intriguing features.
For comparison, we have also studied the scaling of the disorder operator for the square lattice $J_{1}$-$J_{2}$ Heisenberg model with the Hamiltonian
@@
where $\langle ij\rangle$ denotes the $J_{1}$ bond and $\langle ij\rangle^{{}^{\prime}}$ denotes the $J_{2}$ bond. This model has a well-established (2+1)d O(3) QCP at $(J_{2}/J_{1})_{c}=1.90951(1)$ Ma et al. (2018b) separating the Néel phase and a symmetric singlet product phase. As will also shown below, the scaling of disorder operator of this O(3) QCP does not have anomalous log-corrections.